OpenStudy (anonymous):

Simplify negative 8 over 2 ÷ 6 over negative 3. −4 2 –2 4

jimthompson5910 (jim_thompson5910):

When you divide two fractions, you flip the second fraction and multiply \[\Large \frac{-8}{2} \div \frac{6}{-3} = \frac{-8}{2} \times \frac{-3}{6} = ??\]

OpenStudy (anonymous):

-4 x - 1/2?

OpenStudy (anonymous):

hold up i need to eat dinner lol sorry

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

there's no rush

OpenStudy (anonymous):

Ok im back

jimthompson5910 (jim_thompson5910):

hey

jimthompson5910 (jim_thompson5910):

were you able to multiply straight across?

OpenStudy (anonymous):

Well i thought i was only supposed to multiply -8/2 x -3/6

jimthompson5910 (jim_thompson5910):

yes that is correct

jimthompson5910 (jim_thompson5910):

you multiply the numerators together and the denominators together

OpenStudy (anonymous):

ok then after i got -1/8

OpenStudy (anonymous):

24/12 = 2?

jimthompson5910 (jim_thompson5910):

idk about the -1/8 part but yes 24/12 = 2 is correct

OpenStudy (anonymous):

yea i did something wrong there but i fixed it

jimthompson5910 (jim_thompson5910):

I see

OpenStudy (anonymous):

Ok so i have another one i dont know if im right, can you check?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Which fraction is equivalent to fraction with numerator negative 3 and denominator 7? fraction with numerator 3 and denominator negative 7 fraction with numerator negative 3 and denominator negative 7 fraction with numerator negative 7 and denominator 3 fraction with numerator 7 and denominator 3

OpenStudy (anonymous):

i got the first one

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

thank you.

jimthompson5910 (jim_thompson5910):

you're welcome